OpenStudy (anonymous):

how many roots does 4x^2+9x+5

OpenStudy (jackellyn):

4x^2 + 9x + 5 (x+1)(4x+5)

OpenStudy (anonymous):

how many root ?

OpenStudy (jackellyn):

2

OpenStudy (anonymous):

positive or negative ?

OpenStudy (jackellyn):

do you want to know how many there are or what they are? (x+1)(4x+5) x=-1 and x=-5/4

OpenStudy (anonymous):

yeah what they are like wo negative B. one negative and one positive C. two complex D. two positive

OpenStudy (jackellyn):

after you factor a polynomial, you set what you got equal to 0 and solve for x so i just gave the roots to you x=-1 and x=-5/4 you can interpret them.

OpenStudy (anonymous):

thanks (=